OpenStudy (anonymous):

x^2-4/x-8/x-2 anyone know how to divied this

OpenStudy (anonymous):

My math teacher used to say, "Factored polynomials are your friends!" First, you'd factor out x^2-4. That would equal (x-2)(x+2). Then, remember how to deal with these types of fraction? Flip the denominator and multiply. Here's how it would look: \[\frac{ (x-2)(x+2) }{ 1 } * \frac{ (x-2) }{ (x-8) }\] The answer would be \[\frac{ (x-2)^{2} (x+2)}{ (x-8) }\]